Leasing: You are to use your critical thinking skills, collaboration techniques, creative problem solving tools and communication skills under the following scenario:

Your company (Acme Iron) is considering leasing a new computer, you and your team need to perform analysis to support the decision making process. The lease lasts for 9 years. The lease calls for 10 payments of $10,000 per year with the first payment occurring immediately. The computer would cost $70,650 to buy and would be straight-line depreciated to a zero salvage over 9 years. The actual salvage value is negligible because of technological obsolescence. The firm can borrow at a rate of 8%. The corporate tax rate is 30%.

  1. What is the after-tax cash flow from leasing relative to the after-tax cash flow from purchasing in years 1-9?

  1. What is the after-tax cash flow from leasing relative to the after-tax cash flow from purchasing in year 0?

  1. What is the NPV of the lease relative to the purchase?

  1. What would the after-tax cash flow in year 9 be if the asset had a residual value of $500 (ignoring any possible risk differences)?

  1. Do you have a recommendation?

Concept Check: Understanding which cash flows are relevant is key to determining best financing methods or project acceptance. It helps to detail all you assumptions within the model since questions may arise years after the initial construction of the model.

Helpful Hint: Creating a time-line with corresponding cash flows is usually helpful. You should also do the NPV calculations showing your formula so if anyone wishes to change the variables they will know how to proceed.
Answer #1

a). After-tax cash flow from leasing relative to the after-tax cash flow from purchasing in year 1-9 is

after-tax lease payment - after-tax cash flow from purchasing =

lease payment*(1-Tax rate) - (depreciation*Tax rate)

= (-10,000*(1-30%)) - (76,500/9)*30% = -9,550

b). After-tax cash flow from leasing relative to the after-tax cash flow from purchasing in year 0 is

after-tax lease payment - purchase price

= -7,000 -(-76,500) = 69,500

c). NPV of the lease relative to purchase decision = PV of (lease relative to purchase cash flow in year 0) + PV of (lease relative to purchase cash flow from year 1-9)

Discount rate will be after-tax cost of borrowing which will be 8%*(1-30%) = 5.60%

PV of (lease relative to purchase cash flow from year 1-9): PMT = -9,550; N = 9; rate = 5.60%, solve for PV.

PV = -66,102.15

NPV = 69,500 -66,102.15 = 3,397.85

d). If the asset had residual value then the after-tax cash flow in year 9 would be -9,550 (calculated in part b) - after-tax residual value

After-tax residual value = 500*(1-30%) = 350

After-tax cash flow = -9,550 - 350 = -9,900

e). NPV of the incremental cash flows is positive so the asset should be leased.